OpenStudy (anonymous):

8. A force of 80 pounds on a rope is used to pull a box up a ramp inclined at 100 from the horizontal. The rope forms an angle of 330 with the horizontal. How much work is done pulling the box 25 feet along the ramp?

OpenStudy (anonymous):

Here Ramp 10° , rope 33° , with reference on ramp , 33-10 =23° Force on ramp =80 cos 23 =73.64 pounds W= 73.64*25 =1841 p-f